OpenStudy (anonymous):

how do you subtract 4x+2+7x+7+5x-4 - x+3+ x+7+2x-5. i need help ASAP

OpenStudy (anonymous):

4x+2+7x+7+5x-4-x+3+x+7+2x-5 2+11x+7+5x-4-x+3+x+7+2x-5 9+11x+7+5x-4-x+3+7+2x-5 16+11x+5x-4-x+3+7+2x-5 16+16x-4-x+3+7+2x-5 12+16x-x+3+7+2x-5 12+15x+3+7+2x-5 15+15x+7+2x-5 22+15x+2x-5 22+17x-5 17+17x Answer:17+17x

OpenStudy (unklerhaukus):

\[4x+2+7x+7+5x-4-x+3+ x+7+2x-5\]\[=4x+7x+5x-x+x+2x+2+7-4+3+7-5\]\[=(4+7+5-1+1+2)x+2+7-4+3+7-5\]